# End-to-End Framework for Predicting the Remaining Useful Life of Lithium-Ion Batteries

### v2

@@ Abstract
[2]organization=Faculty of Information Technology, Ton Duc Thang University, city=Ho Chi Minh City, postcode=700000, country=Vietnam
[…]
[3]organization=Software Engineering Department, FPT University, city=Da Nang, postcode=550000, country=Vietnam
1 Introduction
The growing global emphasis on sustainable energy has accelerated the advancement of battery technologies in recent years. Consequently, the lithium-ion battery (LIB) market is projected to surpass 170 billion USD by 2030 [28]. LIBs are widely used in applications such as grid energy storage systems, electric vehicles (EVs), and drones, all of which increasingly depend on effective Battery Health Management (BHM) systems [41, 5, 13] to ensure reliable performance. Typical BHM systems include key functionalities such as state of charge (SOC) estimation [32, 39], state of health (SOH) estimation [42, 22, 8], and remaining useful life (RUL) prediction [26, 25]. Among these, RUL prediction—which estimates the number of charge-discharge cycles remaining before a battery reaches its end-of-life (EOL)—is particularly crucial due to its implications for proactive maintenance, and safety assurance. During operation, LIBs undergo multiple forms of degradation during cycling, primarily caused by the loss of lithium inventory (LLI) and loss of active material (LAM) [27]. These challenges underscore the need for accurate RUL prediction models, not only to support real-time diagnostics and timely maintenance, but also to inform battery design [12].
In modern RUL prediction for LIBs, methods are mainly classified into two categories: curve-based and cycle-feature-based approaches. Curve-based methods model the degradation of a health indicator (e.g., capacity) over cycles using predefined functions (e.g., exponential decay), such as the double exponential model (DEM) [23], variational mode decomposition (VMD) [10], or empirical mode decomposition (EMD) [20, 17]. By fitting the historical data, these methods extrapolate the cycle at which the indicator falls below a predefined EOL threshold (e.g., 80% capacity [38]), thereby yielding the predicted number of remaining charge-discharge cycles (RUL). However, these methods have limited flexibility, as they assume a fixed degradation trend throughout the entire battery cycle life, which may not accurately capture real-world variability. Moreover, they often require a long degradation history to make accurate predictions. For instance, the method in [45] relies on data from the first 450 charge-discharge cycles to perform accurate RUL prediction, making it unsuitable for early-stage prognostics. In contrast, cycle-feature-based methods extract meaningful features from a few recent cycles [49, 24], enabling more adaptive, earlier predictions under different operating conditions across charge-discharge cycles.
In recent cycle-feature-based approaches, [49] predicts RUL of LIBs using only features extracted from the current charge-discharge cycle. The features including discharge capacity, initial battery capacity, and working condition-derived variables. However, this approach neglects the short-term degradation trends observable in the preceding cycles, which are critical for accurately capturing the battery’s dynamic aging behavior. Although [49] reports promising results, these outcomes are partially influenced by an imbalanced data split—70 batteries for training and only 7 for testing—which makes the testing set small compared to the training set, resembling a biased setup that may lead to overfitting and reduced generalizability. In another study, [53] employs features extracted from the incremental capacity (IC) curve of a partial segment of the current charge-discharge cycle to predict the SOH and RUL. However, the method is evaluated only on a small battery dataset comprising just four battery cells discharged under constant current. This results in highly similar data distributions between the training and testing sets, making the prediction task easier and less representative of real-world scenarios. In contrast, a more robust and realistic approach is presented in [24], where features are extracted from 10 selected cycles within the 30 most recent charge-discharge cycles to predict RUL. This strategy effectively captures recent degradation patterns, enhancing predictive performance. Furthermore, [24] incorporates transfer learning to leverage degradation knowledge from a different source domain, improving the prediction accuracy in the target domain. Their method is validated on one of the largest publicly available battery datasets, comprising 55 cells for training and 22 cells for testing, with diverse discharge profiles to simulate real-world conditions.
In this work, we adopt the standard approach introduced in [24], where a window of recent sequential charge-discharge cycles is used as input to predict the number of remaining charge-discharge cycles. Our proposed method features a novel signal preprocessing pipeline and a hybrid deep learning (DL) prediction model.
The prediction model comprises two parallel branches. The first branch integrates a 1D Convolutional Neural Network (CNN) block and Attentional Long Short-Term Memory (A-LSTM) network [18] to capture temporal dependencies. The second branch employ an Ordinary Differential Equation-based LSTM (ODE-LSTM) block to enhance the modeling of dynamic sequences. The output from both branches are fused and passed through a fully connected layer to produce a RUL value, representing the predicted number of remaining charge-discharge cycles.
We also evaluate the proposed prediction model using transfer learning under various learning strategies and source data splitting scenarios. The results demonstrate that our model remains robust even when trained on limited source data. The model is validated on two of the most widely used lithium-ion battery datasets: one with varying charging profiles [38] and the other with diverse discharging profiles [24].
The remainder of this paper is organized as follows. Section 2 introduces the proposed method, including signal preprocessing, model architecture, and evaluation metrics. Section 3 describes the datasets and presents data analysis. Section 4 covers the experimental setup, model evaluations, and additional validation with transfer learning. Finally, Section 5 concludes the paper and outlines future work.
2 Proposed Method
The proposed method predicts RUL of lithium-ion batteries, defined as the number of charge-discharge cycles remaining from the current cycle $i$ until the battery reaches its EOL. The approach comprises two main stages: (1) a signal preprocessing pipeline for extracting informative features, and (2) a prediction model based on a hybrid deep learning architecture. The overall architecture is illustrated in Figure 1. The details of these two components are presented in the following subsections. Note that the output from the sigmoid layer is multiplied by a scaling factor of 3000 to obtain the final predicted RUL value. Therefore, the observed RUL labels are divided by the same scaling factor prior to training the model.
Figure 1: Overall architecture of our proposed end-to-end framework.
2.1 Signal Preprocessing
We propose a signal preprocessing pipeline that constructs a feature tensor $\mathcal{F}^{(i)}_{30}$ for each input sample at cycle $i$. The objective of the pipeline is to reduce noise and enhance the salient characteristics of the signals. The proposed preprocessing pipeline consists of the following steps:
  1. 1.
    Capacity Interpolation and Denoising: For each charge-discharge cycle $i$, we first generate evenly spaced values between the minimum and maximum of the current signal $I_{i}(t)$. A one-dimensional interpolation method [29] is then applied, using the sorted current values as the $x$-coordinates and the corresponding discharge capacities as the $y$-coordinates. This yields the derived capacity feature $\dot{Q}_{i}(Q,I)$, a technique we refer to as Capacity Derivative Tracking. Consequently, we obtain the data sequences $I_{i}(t)$, $V_{i}(t)$, $Q_{i}(t)$, and $\dot{Q}_{i}(Q,I)$. Each of these sequences is subsequently denoised using a selected filtering method—specifically, the Savitzky–Golay filter [37]—resulting in the smoothed signals $\tilde{x}_{i}\in\{\tilde{I}_{i}(t),\tilde{V}_{i}(t),\tilde{Q}_{i}(t),\tilde{\dot{Q}}_{i}(Q,I)\}$.
  2. 2.
    Statistical Feature Extraction: From each denoised sequence $\tilde{x}_{i}$, we compute an extracted vector $F$ consisting of six statistical descriptors:
    $F(\tilde{x}_{i})=\left[\mu(\tilde{x}_{i}),\sigma(\tilde{x}_{i}),\min(\tilde{x}_{i}),\max(\tilde{x}_{i}),\mathrm{Var}(\tilde{x}_{i}),\mathrm{Median}(\tilde{x}_{i})\right],$
    where $\mu$ is the mean, $\sigma$ is the standard deviation, $\min$ and $\max$ denote the minimum and maximum values, $\mathrm{Var}$ is the variance, and $\mathrm{Median}$ is the median of the smoothed signal $\tilde{x}_{i}$.
  3. 3.
    Delta Feature Computation: To capture temporal dynamics, we compute the change in derived features $F(\tilde{\dot{Q}}_{i})$ over a cycle interval $\delta$ (set to 9 charge-discharge cycles, based on the optimal configuration identified in Figure 4):
    $\Delta F(\tilde{\dot{Q}}_{i})=F(\tilde{\dot{Q}}_{i})-F(\tilde{\dot{Q}}_{i-\delta}),$
    where $\Delta F(\tilde{\dot{Q}}_{i})$ represents the change in derived features between cycle $i$ and the previous cycle $i-\delta$.
  4. 4.
    Feature Fusion: For each charge-discharge cycle $i$ where $i>\delta$, we extract the following feature vectors: $F(\tilde{I}_{i})$, $F(\tilde{V}_{i})$, $F(\tilde{Q}_{i})$, $\Delta F(\tilde{\dot{Q}}_{i})$, each of size $\mathbb{R}^{6}$. Together, they form a fused feature matrix of size $\mathbb{R}^{4\times 6}$:
    $\mathcal{F}^{(i)}=\left[F(\tilde{I}_{i}),F(\tilde{V}_{i}),F(\tilde{Q}_{i}),\Delta F(\tilde{\dot{Q}}_{i})\right]\in\mathbb{R}^{4\times 6}.$
    For RUL prediction, each input sample comprises data from 10 uniformly spaced cycles selected within a 30-cycle window (i.e., one cycle every three cycles), following the strategy in [24]. The resulting input sample is:
    $\mathcal{F}^{(i)}_{30}\in\mathbb{R}^{10\times 4\times 6}.$
    $\mathcal{F}^{(i)}_{30}$ is reshaped to $\mathbb{R}^{10\times 24}$ for MinMax scaling [31] to the range $(0,1)$ across the 24 dimensions, then reshaped back to $\mathbb{R}^{10\times 4\times 6}$. The scaler is fitted on the training set and applied to the test set.
2.2 Prediction Model
Each input sample $\mathcal{F}^{(i)}_{30}$, as described in the previous section, is fed into the proposed prediction model to estimate the RUL. Our prediction model comprises three primary components: CNN block, A-LSTM block, and ODE-LSTM block. Among these, A-LSTM block is adopted from the prior work [18], while the other components are tailored.
CNN block
The input tensor $\mathcal{F}^{(i)}_{30}$ is first reshaped into a sequence of 10 time steps, each represented by a 24-dimensional feature vector. This sequence is then passed through a stack of three one-dimensional convolutional layers [15] with a shared kernel size of 5 and progressively increasing channel dimensions—$H$, $2H$, and $4H$, where $H=64$—consistent with the optimal configuration identified in Figure 5. Each convolution operation is followed by batch normalization [4], and a Leaky ReLU activation [51]. A dropout [46] layer with a fixed rate of 0.3 is applied at the end of the CNN block for regularization. Note that this fixed dropout rate of 0.3 is consistently used across all dropout layers in the prediction model. The resulting output is a feature map of shape $\mathbb{R}^{10\times 256}$, effectively capturing temporal dependencies and patterns within the cycle window.
A-LSTM Block
The feature map from the CNN block is fed into a two A-LSTM cells [18] with a sequence length of 10. The input size to the A-LSTM is $256$, and the hidden size is $128$. This block captures long-range temporal dependencies with attention mechanisms across the sequence. The hidden state at the final timestep is extracted and regularized via dropout, resulting in a output vector of shape $\mathbb{R}^{128}$.
ODE-LSTM Block
This block processes the input sample $\mathcal{F}^{(i)}_{30}$ by reshaping it into a matrix of shape $\mathbb{R}^{10\times 24}$, while preserving the temporal order. A continuous time sequence $t\in[0,1]$ with 10 uniformly spaced points is generated to simulate continuous-time dynamics. Both the reshaped input and time sequence are fed into two stacked ODE-LSTM cells, with each input size 24 and hidden size 256. The final hidden state is extracted and passed through a dropout layer, resulting in an output vector of size $\mathbb{R}^{256}$.
Our ODE-LSTM cell extends the traditional long short-term memory (LSTM) [40] architecture by integrating neural ordinary differential equations (ODEs) [6] to model the continuous evolution of hidden states between discrete inputs. The computation proceeds in two stages: (1) continuous-time evolution of the hidden state using an ODE solver, and (2) discrete-time update using the LSTM cell, as follows:
$\tilde{h}_{t}$ $=h_{t-1}+\text{ODE}_{\theta}(h_{t-1},t-1,t),$
$i_{t}$ $=\sigma(W_{i}[x_{t};\tilde{h}_{t}]),$
$f_{t}$ $=\sigma(W_{f}[x_{t};\tilde{h}_{t}]),$
$o_{t}$ $=\sigma(W_{o}[x_{t};\tilde{h}_{t}]),$
$\tilde{c}_{t}$ $=\tanh(W_{c}[x_{t};\tilde{h}_{t}]),$
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}_{t},$
$h_{t}$ $=o_{t}\odot\tanh(c_{t}),$
where $\text{ODE}_{\theta}$ models the continuous transition of hidden states and is implemented using Euler integration:
[…]
Here, $f_{\theta}$ is a neural network (implemented as a linear layer), and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes vector concatenation. $x_{t}$ is the input vector at time step $t$. $h_{t}$ is the hidden state at time $t$. $\tilde{h}_{t}$ is the hidden state after ODE evolution from $t_{t-1}$ to $t_{t}$. $c_{t}$ is the cell state at time $t$. $\sigma(\cdot)$ denotes the sigmoid activation function [30], and $\tanh(\cdot)$ denotes the hyperbolic tangent activation [30]. $W_{i}$, $W_{f}$, $W_{o}$, and $W_{c}$ are trainable weight matrices for the input, forget, output, and candidate gates, respectively. $\odot$ denotes element-wise multiplication.
The output features from the CNN + A-LSTM and ODE-LSTM branches are concatenated into a single vector of dimension 384, which is then passed through a linear layer followed by a sigmoid activation function. The resulting value is multiplied by a scaling factor of 3000 to produce the final prediction, representing the RUL. Our proposed hybrid deep learning model is trained using Mean Squared Error (MSE) loss function, optimized with the AdamW optimizer [21]. The MSE loss is defined as:
@@ 2.2 Prediction Model
2.3 Evaluation Metrics
To assess the performance of RUL prediction, we employ three evaluation metrics: Root Mean Square Error ($RMSE$) [47], R-squared ($R^{2}$) [45], and Mean Absolute Percentage Error ($MAPE$) [7]. These metrics are defined as follows:
@@ 2.3 Evaluation Metrics
3 Dataset
3.1 Dataset Description
3.1.1 First dataset
The dataset [38] features a comprehensive study of 124 lithium-ion batteries (LIBs) with LFP/graphite chemistry, focusing on various fast-charging conditions while maintaining consistent discharging parameters. Each battery in the dataset has a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V, with cycle lifespans ranging from 150 to 2,300 cycles, highlighting a broad range of battery longevity. All LIBs underwent uniform discharge procedures, involving a constant current discharge at a rate of 4 C until the voltage reached 2 V, followed by a constant voltage discharge at 2 V until the current dropped to C/50. Charging was conducted at rates between 3.6 C and 6 C under controlled temperature conditions of 30°C in an environmental chamber. The dataset comprises roughly 96,700 cycles, making it one of the largest datasets exploring various fast-charging protocols. It is divided into three sections: a training set with 41 LIBs, a validation set with 43 LIBs, and a test set with 40 LIBs.
3.1.2 Second dataset
The dataset [24], which involved varying discharge conditions while maintaining consistent fast-charging conditions, was collected from a battery degradation experiment. This experiment utilized 77 cells (LFP/graphite A123 APR18650M1A), each with a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. Each cell underwent a distinct multi-stage discharge protocol, while all cells shared the same fast-charging protocol. The experiment was carried out in two thermostatic chambers, with a controlled temperature of 30°C. The dataset contains a total of 146,122 discharge cycles, making it one of the largest datasets that incorporate various discharge protocols. The cycle life of the cells ranges from 1,100 to 2,700 cycles, with an average of 1,898 cycles and a standard deviation of 387 cycles. The discharge capacity as a function of cycle number shows a broad distribution of cycle lives. The dataset is divided into two parts: In 55 LIBs predefined for training, 90% is used for the training set, 10% for the validation set, and the remaining 22 LIBs are used as the test set.
3.2 Data Analysis
@@
The degradation of LIBs through charge-discharge cycles is primarily caused by the LLI, leading to irreversible capacity fade. IC analysis, which plots $dQ/dV$ versus voltage, is a widely used diagnostic tool for identifying LLI-related degradation [1].
@@
As shown in Figure 2, smoothed IC curves during constant current charging illustrate degradation trends in lithium-ion batteries. The position, intensity, and shape of $dQ/dV$ peaks provide key insights into electrochemical aging. Notably, the graphite peak associated with the $\mathrm{LiC_{6}\rightarrow LiC_{12}}$ phase transition [19] shifts left by over 10mV between Cycle 1 and Cycle 1501 (right plot), indicating lithium loss due to solid electrolyte interphase (SEI) formation. This irreversible process consumes lithium ions. Concurrently, the peak height drops by about 35%, signaling a decline in active lithium available for intercalation [33, 50].
@@
Overall, LLI accounts for approximately 70% of the total observed capacity fade, based on the analysis of smoothed IC curves in Figure 2 using the methodology from [33]. Additionally, the figure reveals a capacity reduction of around 20% after 1500 cycles, with the degradation rate varying across the charge-discharge lifespan. These findings highlight the need for a robust RUL prediction method capable of capturing non-uniform aging patterns to reliably predict the RUL of lithium-ion batteries.
4 Experiments and Discussion
4.1 Experimental Setup
@@
Our proposed RUL model is developed using the PyTorch framework. All experiments are performed on an NVIDIA 4090 GPU with 24GB of memory. Each experiment is trained for 10 epochs with a batch size of 128 and a constant learning rate of 0.0005. To minimize variability in the training process, each experiment is repeated 10 times, and the final prediction is calculated as the average of these multiple runs.
4.2 Evaluation of Signal Preprocessing
In the denoising step, to mitigate noise in voltage, current and capacity signals, four denoising techniques including Gaussian smoothing [48], Savitzky-Golay filtering [14], wavelet decomposition [9], and Kalman filtering-along with a no-denoising baseline (raw data), were evaluated on the second dataset. As shown in Figure 3, the Savitzky-Golay filter [36] achieved the optimal performance, yielding an RMSE of 169.93 (cycles). Compared to Gaussian smoothing and wavelet decomposition, the Savitzky-Golay method demonstrates superior capability in peak preservation and adaptive smoothing.
One key factor influencing the effectiveness of the Savitzky-Golay filter is the window length, which balances noise reduction and signal distortion. To determine the optimal configuration, various window lengths were tested. As shown in the right panel of Figure 3, a window length of 191 points provides the best trade-off between denoising and feature preservation. Consequently, the Savitzky-Golay filter with a window length of 191 is adopted in the denoising stage of our signal preprocessing pipeline.
@@ 4.2 Evaluation of Signal Preprocessing
Figure 4 on the left shows that using only derived features results in the highest RMSE of 252.52, indicating that relying solely on derived features leads to poor predictive performance. In contrast, the feature combination of I, V, Q, and dQ(I) achieves the lowest RMSE value of 162.37, outperforming other feature sets. The strong contribution of dQ(I) highlights the effectiveness of the proposed Capacity Derivative Tracking method in estimating discharge data. By aligning discharge capacity values with corresponding current measurements, this method provides a more consistent representation of the current–discharge capacity relationship over time, enhancing the prediction performance of the model.
After the interpolation step, the difference in extracted features between each cycle and the one occurring delta steps earlier is computed. The choice of the delta parameter is important, as it determines the temporal resolution of feature variation. To assess its impact, we evaluated delta values ranging from 1 to 10, as shown in Figure 4 on the right. Among these, a delta value of 9 achieved the best performance, yielding the lowest RMSE of 162.37.
@@
4.3 Validation of Prediction Model
As shown in Table 1, the proposed signal preprocessing pipeline, which includes the derived capacity feature $\dot{Q}_{i}(Q,I)$, consistently improves the performance of all evaluated models. Notably, our proposed prediction model, CNN + A-LSTM + ODE-LSTM, achieves the best results, with the lowest RMSE of 158.09, highest $R^{2}$ score of 0.89, and lowest MAPE of 6.95%. These results confirm the effectiveness of combining CNN with A-LSTM and ODE-LSTM components, especially when enriched with the derived capacity feature.

### v3

@@ II-B Prediction Model
Here, the factor $3000$ scales the sigmoid output from $[0,1]$ to $[0,3000]$ cycles, matching the maximum expected RUL range; thus $\hat{y}$ represents the predicted number of useful remaining cycles (RUL).
Our proposed hybrid deep learning model is trained using Mean Squared Error (MSE) loss to remain consistent with most previous approaches and ensure a fair comparison. In future work, we plan to investigate heteroscedastic loss formulations. The model is optimized with the AdamW optimizer [loshchilov2017decoupled]. The MSE loss is defined as:
@@
II-C Evaluation Metrics
To assess the performance of RUL prediction, we employ three evaluation metrics: Root Mean Square Error ($RMSE$) [willmott2005advantages], R-squared ($R^{2}$) [wang2023remaining], and Mean Absolute Percentage Error ($MAPE$) [de2016mean]. These metrics are defined as follows:
@@ II-C Evaluation Metrics
III Dataset
III-A Dataset Description
First dataset
The dataset [severson2019data] features a comprehensive study of 124 lithium-ion batteries (LIBs) with LFP/graphite chemistry, focusing on various fast-charging conditions while maintaining consistent discharging parameters. Each battery in the dataset has a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V, with cycle lifespans ranging from 150 to 2,300 cycles, highlighting a broad range of battery longevity. All LIBs underwent uniform discharge procedures, involving a constant current discharge at a rate of 4 C until the voltage reached 2 V, followed by a constant voltage discharge at 2 V until the current dropped to C/50. Charging was conducted at rates between 3.6 C and 6 C under controlled temperature conditions of 30°C in an environmental chamber. The dataset comprises roughly 96,700 cycles, making it one of the largest datasets exploring various fast-charging protocols.
In this dataset, battery EOL is defined at 80% of the nominal capacity (i.e., 0.88 Ah for these cells). While the dataset provides extensive coverage across fast-charging policies, it is imbalanced with respect to cycle life: many cells reach EOL relatively early (150–600 cycles), whereas only a smaller subset survives beyond 2,000 cycles. This imbalance poses challenges for model generalization, as predictive accuracy can be biased toward short-lived cells.
Following the official split reported in the Supplementary Information of [severson2019data], we use 41 LIBs for training, 43 for validation, and 40 for testing. The training and validation sets are drawn from the same manufacturing batches (2017-05-12 and 2017-06-30) and the same charging-policy families, but are disjoint by cell. For most policies, one LIB is assigned to the training set and its matched counterpart to the validation set, yielding an in-distribution holdout of unseen cells. The test set comprises all LIBs from the later 2018-04-12 batch, serving as an out-of-distribution holdout due to batch and protocol shift.
@@ III-A Dataset Description
Second dataset
The dataset [ma2022real], which involved varying discharge conditions while maintaining consistent fast-charging conditions, was collected from a battery degradation experiment. This experiment utilized 77 cells (LFP/graphite A123 APR18650M1A), each with a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. Each cell underwent a distinct multi-stage discharge protocol, while all cells shared the same fast-charging protocol. The experiment was carried out in two thermostatic chambers, with a controlled temperature of 30°C. The dataset contains a total of 146,122 discharge cycles, making it one of the largest datasets that incorporate various discharge protocols. The cycle life of the cells ranges from 1,100 to 2,700 cycles, with an average of 1,898 cycles and a standard deviation of 387 cycles. The discharge capacity as a function of cycle number shows a broad distribution of cycle lives.
In this dataset, EOL is also defined at 80% of the nominal capacity (0.88 Ah). Compared to the first dataset, the discharge-protocol variability introduces additional heterogeneity in cycle lengths and degradation trajectories. While many cells operate stably over long periods (surviving more than 2,500 cycles), others reach EOL earlier due to aggressive discharge protocols. This imbalance in cycle life distributions challenges the model to learn from both long-lived and short-lived cells.
Following the official split in [ma2022real], the dataset is divided into two parts: In 55 LIBs predefined for training, 90% is used for the training set, 10% for the validation set, and the remaining 22 LIBs are used as the test set. This protocol-disjoint split intentionally induces a distribution shift between training and testing: because each set contains different discharge protocols, their current profiles, voltage–capacity trajectories, cycle lengths, and operating patterns differ. It provides a stricter evaluation of robustness and transferability to unseen operating conditions.
In both datasets, careful control of operating conditions ensures that degradation patterns is attributed primarily to the designed experimental protocols rather than uncontrolled external factors. In the first dataset, charging rate and temperature were tightly regulated, resulting in 96,700 cycles and making it one of the largest publicly available resources for studying the interplay between fast-charging strategies and battery aging. In the second dataset, discharge rate and temperature were carefully controlled, yielding 146,122 cycles and providing one of the largest resources for investigating the interplay between discharge strategies and battery aging.
III-B Data Analysis
@@
The degradation of LIBs through charge-discharge cycles is primarily caused by the LLI, leading to irreversible capacity fade. IC analysis, which plots $dQ/dV$ versus voltage, is a widely used diagnostic tool for identifying LLI-related degradation [ansean2019lithium].
@@
As shown in Figure 2, smoothed IC curves during constant current charging illustrate degradation trends in lithium-ion batteries. The position, intensity, and shape of $dQ/dV$ peaks provide key insights into electrochemical aging. Notably, the graphite peak associated with the $\mathrm{LiC_{6}\rightarrow LiC_{12}}$ phase transition [Kinetically] shifts left by over 10mV between Cycle 1 and Cycle 1501 (right plot), indicating lithium loss due to solid electrolyte interphase (SEI) formation. This irreversible process consumes lithium ions. Concurrently, the peak height drops by about 35%, signaling a decline in active lithium available for intercalation [Pinson_2013, 7488267].
@@
Overall, LLI accounts for approximately 70% of the total observed capacity fade, based on the analysis of smoothed IC curves in Figure 2 using the methodology from [Pinson_2013]. Additionally, the figure reveals a capacity reduction of around 20% after 1500 cycles, with the degradation rate varying across the charge-discharge lifespan. These findings highlight the need for a robust RUL prediction method capable of capturing non-uniform aging patterns to reliably predict the RUL of lithium-ion batteries.
IV Experiments and Discussion
IV-A Experimental Setup
@@
Our proposed RUL model is developed using the PyTorch framework. All experiments are performed on an NVIDIA 4090 GPU with 24GB of memory. Each experiment is trained for 10 epochs with a batch size of 128 and a constant learning rate of 0.0005. We train the model for 10 epochs, as the compact feature representation produced by our signal preprocessing pipeline allows the network to converge rapidly. Each input signal (e.g., current, voltage) is compressed into six representative points, rather than 100 resamples [ma2022real] or dozens of handcrafted features [severson2019data], which significantly reduces the model’s computational complexity while preserving essential degradation information. We also tested learning rates of 0.001, 0.0005, 0.0002, and 0.0001. 0.0005 consistently provided the best validation and test performance with the most stable training. To minimize variability in the training process, each experiment is repeated 10 times, and the final prediction is calculated as the average of these multiple runs.
@@
IV-B Evaluation of Signal Preprocessing
In the denoising step, to mitigate noise in voltage, current and capacity signals, four denoising techniques including Gaussian smoothing [wink2004denoising], Savitzky-Golay filtering [karaim2019low], wavelet decomposition [ergen2012signal], and Kalman filtering-along with a no-denoising baseline (raw data), were evaluated on the second dataset. As shown in Figure 3, the Savitzky-Golay filter [sayadi2008ecg] achieved the optimal performance, yielding an RMSE of 169.93 (cycles). Compared to Gaussian smoothing and wavelet decomposition, the Savitzky-Golay method demonstrates superior capability in peak preservation and adaptive smoothing.
One key factor influencing the effectiveness of the Savitzky–Golay filter is the window length, which governs the balance between noise reduction and signal distortion. To determine the optimal configuration, various window lengths were evaluated. As shown in the right panel of Figure 3, a window length of 191 points yields the lowest RMSE value, whereas both shorter (181) and longer (199) windows resulted in higher errors. A smaller window length insufficiently smooths the raw signals, leaving residual measurement noise that propagates into feature extraction and model training. Conversely, a larger window over-smooths the signal, attenuating transient behaviors such as end-of-discharge voltage drops. The chosen window length of 191 thus achieves an optimal balance between noise suppression and feature fidelity, preserving the physical integrity that is essential for accurate battery RUL prediction. Consequently, the Savitzky–Golay filter with a window length of 191 is adopted in the denoising stage of our signal preprocessing pipeline.